# Axi-majoron for almost everything

### v2

@@ Acknowledgments
This work was supported by the National Research Foundation of Korea grants by the Korea government: 2017R1D1A1B06035959 (W.I.P.), 2022R1A4A5030362 (W.I.P.), NRF-2019R1A2C3005009 (P.K.) and KIAS Individual Grants under Grant No. PG021403 (P.K.). It was also supported by the Spanish grants CIPROM/2021/054 (Generalitat Valenciana), PID2020-113775GB-I00 (AEI/10.13039/501100011033) (G.B. and W.I.P.) and by the European ITN project HIDDeN (H2020-MSCA-ITN-2019/860881-HIDDeN) (G.B.).
Appendix A Metric component in a various basis
@@
$\frac{\partial\Omega}{\partial\phi}$ $=$ $\frac{2\xi_{\phi}}{M_{\rm P}^{2}}\phi\left[1+\alpha\cos\left(2\theta_{\Phi}\right)\right]\ ,$ (A.9)
$\frac{\partial\Omega}{\phi\partial\theta}$ $=$ $-\frac{2\xi_{\phi}}{M_{\rm P}^{2}}\phi\alpha\sin\left(2\theta_{\Phi}\right)\ ,$ (A.10)
@@
$K_{22}$ $=$ $\Omega+6\left(\frac{\xi_{\phi}\phi}{M_{\rm P}}\right)^{2}\left[\alpha\sin\left(2\theta_{\Phi}\right)\right]^{2}\ ,$ (A.12)
$K_{12}$ $=$ $-6\left(\frac{\xi_{\phi}\phi}{M_{\rm P}}\right)^{2}\left[1+\alpha\cos\left(2\theta_{\Phi}\right)\right]\left[\alpha\sin\left(2\theta_{\Phi}\right)\right]\ .$ (A.13)
Appendix B Derivatives of the Einstein frame potential w.r.t. the canonical fields
In the large field region of $\xi_{\phi}^{1/2}\phi\gg M_{\rm P}$, in terms of the approximate canonical fields, the potential is read as
@@
Let us define $A_{1}\equiv 2\sqrt{\xi_{\phi}/(1+6\xi_{\phi})}$ and $A_{2}\equiv 2\sqrt{\xi_{\phi}}$ for notational convenience. Then, derivatives of the potential are found to be
$\frac{\partial U}{\partial\varphi_{1}}$ $=$ $\frac{2A_{1}U/M_{\rm P}}{\left[1+\left(1+\alpha\cos\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)\right)\exp\left[\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\right]}$ (B.2)
$\frac{\partial^{2}U}{\partial\varphi_{1}^{2}}$ $=$ $\frac{2A_{1}^{2}\left[2-\left(1+\alpha\cos\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)\right)\exp\left[\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\right]U/M_{\rm P}^{2}}{\left[1+\left(1+\alpha\cos\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)\right)\exp\left[\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\right]^{2}}$ (B.3)
$\frac{\partial^{3}U}{\partial\varphi_{1}^{3}}$ $=$ $\frac{2A_{1}^{3}\left[4-7\exp\left[\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\left(1+\alpha\cos\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)\right)+\exp\left[\frac{2A_{1}\varphi_{1}}{M_{\rm P}}\right]\left(1+\alpha\cos\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)\right)^{2}\right]}{\left(U/M_{\rm P}^{3}\right)^{-1}\left[1+\left(1+\alpha\cos\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)\right)\exp\left[\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\right]^{3}}$ (B.4)
$\frac{\partial U}{\partial\varphi_{2}}$ $=$ $\frac{2A_{2}\alpha\sin\left(A_{2}\varphi_{2}/M_{\rm P}\right)U/M_{\rm P}}{\left[1+\alpha\cos\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)+\exp\left[-\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\right]}$ (B.5)
$\frac{\partial^{2}U}{\partial\varphi_{2}^{2}}$ $=$ $\frac{2A_{2}^{2}\alpha\exp\left[\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\left[\left(1+\left(1-\alpha\right)\exp\left[\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\right)\cos\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)+2\alpha\exp\left[\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\right]}{\left(U/M_{\rm P}^{2}\right)^{-1}\left[1+\left(1+\alpha\cos\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)\right)\exp\left[\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\right]^{2}}$ (B.6)
$\frac{\partial^{3}U}{\partial\varphi_{2}^{3}}$ $=$ $\frac{2A_{2}^{3}\alpha\exp\left[\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]f(\varphi_{1},\varphi_{2})\sin\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)}{\left(U/M_{\rm P}^{3}\right)^{-1}\left[1+\left(1+\alpha\cos\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)\right)\exp\left[\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\right]^{3}}$ (B.7)
$\frac{\partial^{2}U}{\partial\varphi_{1}\partial\varphi_{2}}$ $=$ $\frac{6A_{1}A_{2}\alpha\exp\left[\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\sin\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)U/M_{\rm P}^{2}}{\left[1+\left(1+\alpha\cos\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)\right)\exp\left[\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\right]^{2}}\ ,$ (B.8)
where
$f(\varphi_{1},\varphi_{2})$$\equiv$ $-1+\exp\left[\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\left[-2+7\alpha\cos\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)\right]$ (B.9)
$+\exp\left[\frac{2A_{1}\varphi_{1}}{M_{\rm P}}\right]\left[-1+10\alpha^{2}+7\alpha\cos\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)-2\alpha^{2}\cos\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)\right]\ .$
For $\alpha\lll 1$ they are approximated as
$\frac{\partial U}{\partial\varphi_{1}}$ $=$ $2A_{1}\exp\left[-\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\frac{U}{M_{\rm P}}$ (B.10)
$\frac{\partial^{2}U}{\partial\varphi_{1}^{2}}$ $=$ $-2A_{1}^{2}\exp\left[-\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\frac{U}{M_{\rm P}^{2}}$ (B.11)
$\frac{\partial^{3}U}{\partial\varphi_{1}^{3}}$ $=$ $2A_{1}^{3}\left(-7+\exp\left[\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\right)\exp\left[-2\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\frac{U}{M_{\rm P}^{3}}$ (B.12)
$\frac{\partial U}{\partial\varphi_{2}}$ $=$ $2A_{2}\alpha\sin\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)\frac{U}{M_{\rm P}}$ (B.13)
$\frac{\partial^{2}U}{\partial\varphi_{2}^{2}}$ $=$ $2A_{2}^{2}\alpha\cos\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)\frac{U}{M_{\rm P}^{2}}$ (B.14)
$\frac{\partial^{3}U}{\partial\varphi_{2}^{3}}$ $=$ $-2A_{2}^{3}\alpha\sin\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)\frac{U}{M_{\rm P}^{3}}\xrightarrow[]{\theta_{\Phi}\ll 1}-4\alpha\theta_{\Phi}\frac{\lambda_{\Phi}M_{\rm P}}{\sqrt{\xi_{\phi}}}=-8\sqrt{3/\lambda_{\Phi}}\alpha\theta_{\Phi}H_{I}\ .$ (B.15)
@@ C.4 $N+N\rightarrow N+N$
Appendix D Decay rate of right-handed neutrinos
For reader’s convenience, here we collect the partial decay rates of RHNs to 2-body final states from Ref. [78, 79].
  • $m_{N}>m_{W}$: The decay rate is given by [78]
    $\Gamma^{N}=\frac{y_{\nu}^{2}m_{N}}{64\pi}\times\sum_{i}n_{i}\Theta(m_{N}-m_{i})\left(1-\frac{m_{i}^{2}}{m_{N}^{2}}\right)^{2}\left(1+\frac{2m_{i}^{2}}{m_{N}^{2}}\right)^{n_{V,i}}$ (D.1)
    where $i=(h,Z,W)$, $n_{i}=(1,1,2)$ is the number of channels, $n_{V,i}=(0,1,1)$ is the vector boson number, and $\Theta$ is the Heaviside step function.
  • $m_{N}<m_{W}$: The total 2-body rate of $N$ is given by [79]
    $\Gamma^{N}=\frac{|U_{\alpha}|^{2}}{32\pi}G_{F}^{2}m_{N,0}^{3}\times\sum_{i}\mathcal{F}_{i}$ (D.2)
@@ Appendix D Decay rate of right-handed neutrinos
    $|U_{\alpha}|^{2}=\frac{y_{\nu}^{2}v_{h}^{2}}{4m_{N}^{2}}=\frac{m_{\nu}}{2m_{N}}$ (D.3)
    and $\mathcal{F}_{i}$ for each decay channel is given by
    $\pi^{0}+\nu_{\alpha}$ ::: $f_{\pi}^{2}\left(1-\frac{m_{\pi}^{2}}{m_{N}^{2}}\right)^{2}$ (D.4)
    $H^{+}\ell_{\alpha}^{-}$ ::: $2f_{H}^{2}|V_{H}|^{2}\left[\left(1-\frac{m_{\ell}^{2}}{m_{N}^{2}}\right)^{2}-\frac{m_{H}^{2}}{m_{N}^{2}}\left(1+\frac{m_{\ell}^{2}}{m_{N}^{2}}\right)\right]$ (D.5)
    ×$\times\sqrt{\left(1-\frac{(m_{H}-m_{\ell})^{2}}{m_{N}^{2}}\right)\left(1-\frac{(m_{H}+m_{\ell})^{2}}{m_{N}^{2}}\right)}$
    $\eta+\nu_{\alpha}$ ::: $f_{\eta}^{2}\left(1-\frac{m_{\eta}^{2}}{m_{N}^{2}}\right)^{2}$ (D.6)
    $\eta^{\prime}+\nu_{\alpha}$ ::: $f_{\eta^{\prime}}^{2}\left(1-\frac{m_{\eta^{\prime}}^{2}}{m_{N}^{2}}\right)^{2}$ (D.7)
    $\rho^{+}\ell_{\alpha}^{-}$ ::: $4\frac{g_{\rho}^{2}}{m_{\rho}^{2}}|V_{ud}|^{2}\left[\left(1-\frac{m_{\ell}^{2}}{m_{N}^{2}}\right)^{2}-\frac{m_{\rho}^{2}}{m_{N}^{2}}\left(1+\frac{m_{\ell}^{2}-2m_{\rho}^{2}}{m_{N}^{2}}\right)\right]$ (D.8)
    ×$\times\sqrt{\left(1-\frac{(m_{\rho}-m_{\ell})^{2}}{m_{N}^{2}}\right)\left(1-\frac{(m_{\rho}+m_{\ell})^{2}}{m_{N}^{2}}\right)}$
    $\rho^{0}+\nu_{\alpha}$ ::: $2\frac{g_{\rho}^{2}}{m_{\rho}^{2}}\left(1-\frac{m_{\rho}^{2}}{m_{N}^{2}}\right)^{2}\left(1+\frac{2m_{\rho}^{2}}{m_{N}^{2}}\right)$ (D.9)
    with
    $f_{\pi}=130{\rm MeV},f_{\eta}=1.2f_{\pi},\ f_{\eta^{\prime}}=-0.45f_{\pi},\ g_{\rho}=0.102{\rm GeV}^{2},$ (D.10)
    and $f_{H}$ and $V_{H}$ are the decay constants and the CKM matrix associated with the final state mesons, $H^{+}=\pi^{+},K^{+},D^{+},D_{s}^{+},B^{+},B_{c}^{+}$ [79].
Fig 13 shows the total decay rate, $\Gamma_{N_{i}}$, of the mass-eigenstate $N_{i}$ as a function of the mass, $m_{N_{i}}$, where $y_{\nu}$ was assumed to have the seesaw relation in Eq. (3.57).
Figure 13: Decay rate of a RHN mass-eigenstate as a function of its mass $m_{N}$. The vertical dotted line corresponds to $m_{N}=m_{W}$. The horizontal line is a reference line of the expansion rate at the electroweak scale background temperature in the standard radiation-domination universe. It is assumed that RHN decay is flavor-blind for left-handed neutrinos and charged leptons separately. Hence, for a neutrino channel and lepton channel a factor three was multiplied, respectively.
Appendix E The helicity asymmetry of $N_{1}$s at the epoch of their decays
$Y_{1,\rm d}\equiv\frac{\Delta n_{1,\rm d}}{s_{\rm d}}$ (E.1)
E.1 Case of $\phi_{*}<\phi_{\times}$
In this case, one finds
$\Delta n_{1,\rm d}=\Delta n_{1,\times}\left(\frac{a_{\times}}{a_{\rm d}}\right)^{3}$ (E.2)
$s_{\rm d}=\frac{4}{3}\frac{\rho_{1,\rm d}}{T_{\rm d}}$ (E.3)
where $\rho_{1,\rm d}$ is the energy density of $N_{1}$ at its decay. One finds
$\rho_{1,\rm d}={\rm B}_{1}\rho_{\Phi,\times}\left(\frac{a_{*}}{a_{1,\rm eq}}\right)\left(\frac{a_{\times}}{a_{\rm d}}\right)^{3}$ (E.4)
where the energy density of $\Phi$ when $\phi=\phi_{\times}$ is given by
$\rho_{\Phi,\times}=\frac{\lambda_{\phi}}{4}\phi_{0}^{4}=\frac{1}{4}\times\frac{\lambda_{\phi}}{4}\phi_{\times}^{4}=\frac{1}{4}\times\frac{3}{4}T_{\Phi,\times}s_{\Phi,\times}$ (E.5)
with $T_{\Phi,\times}$ defined as
$T_{\Phi,\times}=\left(\frac{\pi^{2}}{30}g_{*}^{\Phi}\right)^{-1/4}\left(\frac{\lambda_{\phi}}{4}\phi_{\times}^{4}\right)^{1/4}\ .$ (E.6)
As a result, one finds
$Y_{1,\rm d}=4\Delta_{\rm AD}\left(\frac{{\tilde{H}}_{*}}{{\tilde{H}}_{1,\rm eq}}\right)^{1/2}\frac{T_{\rm d}}{T_{\Phi,\times}}Y_{\Phi,e}\ .$ (E.7)
E.2 Case of $\phi_{*}>\phi_{\times}$
In this case, one finds
$\Delta n_{1,\rm d}=\Delta n_{1,*}\left(\frac{a_{*}}{a_{\rm d}}\right)^{3}$ (E.8)
$\rho_{1,\rm d}={\rm B}_{1}\rho_{\Phi,*}\left(\frac{a_{*}}{a_{1,\rm eq}}\right)\left(\frac{a_{*}}{a_{\rm d}}\right)^{3}$ (E.9)
where
$\rho_{\Phi,*}=\frac{\lambda_{\phi}}{4}\phi_{*}^{4}=\frac{3}{4}T_{\Phi,*}s_{\Phi,*}$ (E.10)
with $T_{\Phi,*}$ defined in the same way as Eq. (E.6) but $\phi_{\times}$ replaced to $\phi_{*}$. As a result, one finds
$Y_{1,\rm d}=\left(\frac{H_{*}}{H_{1,\rm eq}}\right)^{1/2}\frac{T_{\rm d}}{T_{\Phi,*}}Y_{\Phi,e}\ .$ (E.11)

### v3

@@ Acknowledgments
This work was supported by the National Research Foundation of Korea grants by the Korea government: No. 2017R1D1A1B06035959 (W.I.P.), No. 2022R1A4A5030362 (W.I.P.), No. NRF-2019R1A2C3005009 (P.K.) and KIAS Individual Grants under Grant No. PG021403 (P.K.). It was also supported by the Spanish grants No. CIPROM/2021/054 (Generalitat Valenciana), No. PID2020-113775GB-I00 (AEI/10.13039/501100011033) (G.B. and W.I.P.) and by the European ITN project HIDDeN (No. H2020-MSCA-ITN-2019/860881-HIDDeN) (G.B.).
Appendix A Metric component in various bases
@@
$\frac{\partial\Omega}{\partial\phi}$ $=$ $\frac{2\xi_{\phi}\phi}{M_{\rm P}^{2}}\left[1+\alpha\cos\left(2\theta_{\Phi}\right)\right]\ ,$ (A.9)
$\frac{\partial\Omega}{\partial\theta}$ $=$ $-\frac{2\xi_{\phi}\phi^{2}}{M_{\rm P}^{2}}\alpha\sin\left(2\theta_{\Phi}\right)\ ,$ (A.10)
@@
$K_{22}$ $=$ $\Omega\phi^{2}+6\left(\frac{\xi_{\phi}\phi}{M_{\rm P}}\right)^{2}\phi^{2}\left[\alpha\sin\left(2\theta_{\Phi}\right)\right]^{2}\ ,$ (A.12)
$K_{12}$ $=$ $-6\left(\frac{\xi_{\phi}\phi}{M_{\rm P}}\right)^{2}\phi\left[1+\alpha\cos\left(2\theta_{\Phi}\right)\right]\left[\alpha\sin\left(2\theta_{\Phi}\right)\right]\ .$ (A.13)
Appendix B Derivatives of the Einstein frame potential with respect to the canonical fields
In the large field region of $\xi_{\phi}^{1/2}\phi\gg M_{\rm P}$, in terms of the approximate canonical fields of the polar field-space coordinates, the potential is read as
@@ Appendix B Derivatives of the Einstein frame potential with respect to the canonical fields
Let us define $A_{1}\equiv 2\sqrt{\xi_{\phi}/(1+6\xi_{\phi})}$ and $A_{2}\equiv 2\sqrt{\xi_{\phi}}$ for notational convenience. Then, derivatives of the potential up to the second order are found to be
$\frac{\partial U}{\partial\varphi_{1}}$ $=$ $\frac{2A_{1}U/M_{\rm P}}{\left[1+\left(1+\alpha\cos\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)\right)\exp\left[\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\right]}\ ,$ (B.2)
$\frac{\partial^{2}U}{\partial\varphi_{1}^{2}}$ $=$ $\frac{2A_{1}^{2}\left[2-\left(1+\alpha\cos\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)\right)\exp\left[\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\right]U/M_{\rm P}^{2}}{\left[1+\left(1+\alpha\cos\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)\right)\exp\left[\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\right]^{2}}\ ,$ (B.3)
$\frac{\partial U}{\partial\varphi_{2}}$ $=$ $\frac{2A_{2}\alpha\sin\left(A_{2}\varphi_{2}/M_{\rm P}\right)U/M_{\rm P}}{\left[1+\alpha\cos\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)+\exp\left[-\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\right]}\ ,$ (B.4)
$\frac{\partial^{2}U}{\partial\varphi_{2}^{2}}$ $=$ $\frac{2A_{2}^{2}\alpha\exp\left[\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\left[\left(1+\left(1-\alpha\right)\exp\left[\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\right)\cos\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)+2\alpha\exp\left[\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\right]}{\left(U/M_{\rm P}^{2}\right)^{-1}\left[1+\left(1+\alpha\cos\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)\right)\exp\left[\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\right]^{2}}\ ,$ (B.5)
$\frac{\partial^{2}U}{\partial\varphi_{1}\partial\varphi_{2}}$ $=$ $\frac{6A_{1}A_{2}\alpha\exp\left[\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\sin\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)U/M_{\rm P}^{2}}{\left[1+\left(1+\alpha\cos\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)\right)\exp\left[\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\right]^{2}}\ .$ (B.6)
For $\alpha\lll 1$, they are approximated as
$\frac{\partial U}{\partial\varphi_{1}}$ $=$ $2A_{1}\exp\left[-\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\frac{U}{M_{\rm P}}\ ,$ (B.7)
$\frac{\partial^{2}U}{\partial\varphi_{1}^{2}}$ $=$ $-2A_{1}^{2}\exp\left[-\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\frac{U}{M_{\rm P}^{2}}\ ,$ (B.8)
$\frac{\partial U}{\partial\varphi_{2}}$ $=$ $2A_{2}\alpha\sin\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)\frac{U}{M_{\rm P}}\ ,$ (B.9)
$\frac{\partial^{2}U}{\partial\varphi_{2}^{2}}$ $=$ $2A_{2}^{2}\alpha\cos\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)\frac{U}{M_{\rm P}^{2}}\ ,$ (B.10)
$\frac{\partial^{2}U}{\partial\varphi_{1}\partial\varphi_{2}}$ $=$ $6A_{1}A_{2}\alpha\exp\left[-\frac{A_{1}\varphi_{1}}{M_{\rm P}}\right]\sin\left(\frac{A_{2}\varphi_{2}}{M_{\rm P}}\right)\frac{U}{M_{\rm P}^{2}}\ .$ (B.11)
@@ C.4 $N+N\rightarrow N+N$
Appendix D The helicity asymmetry of $N_{1}$s at the epoch of their decays
$Y_{1,\rm d}\equiv\frac{\Delta n_{1,\rm d}}{s_{\rm d}}$ (D.1)
D.1 Case of $\phi_{*}<\phi_{\times}$
In this case, one finds
$\Delta n_{1,\rm d}=\Delta n_{1,\times}\left(\frac{a_{\times}}{a_{\rm d}}\right)^{3}$ (D.2)
$s_{\rm d}=\frac{4}{3}\frac{\rho_{1,\rm d}}{T_{\rm d}}$ (D.3)
where $\rho_{1,\rm d}$ is the energy density of $N_{1}$ at its decay. One finds
$\rho_{1,\rm d}={\rm B}_{1}\rho_{\Phi,\times}\left(\frac{a_{*}}{a_{1,\rm eq}}\right)\left(\frac{a_{\times}}{a_{\rm d}}\right)^{3}$ (D.4)
where the energy density of $\Phi$ when $\phi=\phi_{\times}$ is given by
$\rho_{\Phi,\times}=\frac{\lambda_{\phi}}{4}\phi_{0}^{4}=\frac{1}{4}\times\frac{\lambda_{\phi}}{4}\phi_{\times}^{4}=\frac{1}{4}\times\frac{3}{4}T_{\Phi,\times}s_{\Phi,\times}$ (D.5)
with $T_{\Phi,\times}$ defined as
$T_{\Phi,\times}=\left(\frac{\pi^{2}}{30}g_{*}^{\Phi}\right)^{-1/4}\left(\frac{\lambda_{\phi}}{4}\phi_{\times}^{4}\right)^{1/4}\ .$ (D.6)
As a result, one finds
$Y_{1,\rm d}=4\Delta_{\rm AD}\left(\frac{{\tilde{H}}_{*}}{{\tilde{H}}_{1,\rm eq}}\right)^{1/2}\frac{T_{\rm d}}{T_{\Phi,\times}}Y_{\Phi,e}\ .$ (D.7)
D.2 Case of $\phi_{*}>\phi_{\times}$
In this case, one finds
$\Delta n_{1,\rm d}=\Delta n_{1,*}\left(\frac{a_{*}}{a_{\rm d}}\right)^{3}$ (D.8)
$\rho_{1,\rm d}={\rm B}_{1}\rho_{\Phi,*}\left(\frac{a_{*}}{a_{1,\rm eq}}\right)\left(\frac{a_{*}}{a_{\rm d}}\right)^{3}$ (D.9)
@@ D.2 Case of $\phi_{*}>\phi_{\times}$
$\rho_{\Phi,*}=\frac{\lambda_{\phi}}{4}\phi_{*}^{4}=\frac{3}{4}T_{\Phi,*}s_{\Phi,*}$ (D.10)
with $T_{\Phi,*}$ defined in the same way as Eq. (D.6) but $\phi_{\times}$ replaced to $\phi_{*}$. As a result, one finds
$Y_{1,\rm d}=\left(\frac{H_{*}}{H_{1,\rm eq}}\right)^{1/2}\frac{T_{\rm d}}{T_{\Phi,*}}Y_{\Phi,e}\ .$ (D.11)